OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

\[\cos \Theta = \frac{ 6+48 }{ \sqrt{20}\sqrt{75} }\]

OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

\[\cos \Theta = \frac{ 54 }{ 10\sqrt{15} }\]

OpenStudy (anonymous):

it only gives me degrees: A) 3.0° B) 6.0° C) -7.0° D)16.0° As answers and i don't know how to change all of that into a degree